OpenStudy (anonymous):

how do you find 15% of 2,450?!?!

OpenStudy (ja1):

you divide the percent by the number and multiply the product by 100

OpenStudy (ja1):

\[P = \frac{ 15 }{ 2450 }*100\]

OpenStudy (anonymous):

ohhh okay!

OpenStudy (uri):

noo. \[\frac{ 15 }{ 100 } \times 2450\]

OpenStudy (ja1):

*Facepalm* she's right and I need some sleep >_<

OpenStudy (uri):

lulz.:3 me too.